How many moles of H2S gas are contained in a 77.8-g sample of this gas?
1
Expert's answer
2020-07-07T14:30:22-0400

Solution.

M(H2S)=34 g/mol

n(H2S)=m(H2S)/M(H2S)=77.8/34=2.29 mol

Answer: 2.29 mol